OpenStudy (anonymous):

Solve: x^2− 2x − 24 = 0. A. x = 4, x = 6 B. x = 4, x = −6 C. x = −4, x = 6 D. x = −4, x = −6

Parth (parthkohli):

Use the quadratic formula to solve.

Parth (parthkohli):

\(\Large \color{Black}{\Rightarrow x = {-b \pm \sqrt{b^2 - 4ac} \over 2a} }\)

OpenStudy (anonymous):

Can't you also factor and then solve? So you would get (x-6)(x+4)? And then you set them both equal to zero like this? (x-6)=0 and (x+4)=0, so your answers would be x=6 and x=-4??

Parth (parthkohli):

Eeky @iheartfood

Parth (parthkohli):

You shouldn't be giving the complete answer

OpenStudy (anonymous):

What do you mean by eeky? hahaa oh oops my bad.. should i delete that post then??

OpenStudy (anonymous):

@ParthKohli should i delete the answer then?

Parth (parthkohli):

No, too late now :P

OpenStudy (anonymous):

aww man!! sorry!! i didn't' know!! my bad...... but was my way okay too @ParthKohli ??

Parth (parthkohli):

Yeah, factoring is a very good way to solve ^_^

OpenStudy (anonymous):

yay!! :)

Parth (parthkohli):

I used to solve them by factoring, but now I see that the formula is convenient.